# Fine-grained Abnormality Prompt Learning for Zero-shot Anomaly Detection

### v1

@@ 3.4 Learning to Select Data-dependent Abnormality Prior
One issue in ZSAD is that the fine-grained abnormality patterns can be very different from the auxiliary dataset to test datasets. In addition to the learning of a set of complementary fine-grained abnormality prompts, it is important to ensure that the learned fine-grained abnormality patterns are generalized to target testing datasets. Inspired by the instance-conditional information design in CoCoOp (Zhou et al., 2022a), we introduce the DAP module to enhance the cross-dataset generalizability of the abnormal tokens in CAP by adaptively selecting the embeddings of the most abnormal regions to serve as a sample-wise abnormality prior for each image input. Particularly, given a query/test image $x$, DAP selects the most abnormal image patches as the abnormality prior to be fed into CAP for assisting the abnormality prompt learning. It achieves this by picking the top $M$ patches whose token embeddings are most similar to the abnormality prompt prototype $\mathbf{F}_{a}$:
@@
where $[\cdot]^{\intercal}$ denotes a transpose operation, $\mathbf{F}_{v}(i,j)$ is the token embedding of the patch centered at $(i,j)$ and $\mathbf{S}_{x}^{a}(i,j)$ is a patch-level anomaly score. The corresponding normal scores can be calculated via $\mathbf{S}_{x}^{n}(i,j)$ using the similarity to $\mathbf{F}_{n}$ in the numerator in Eq. 3.
Let $\textbf{p}_{x}=\{p_{1},p_{2},...p_{M}\}$ be the top $M$ patch embeddings of $x$, FAPrompt then adds additional learnable layers $\psi(\cdot)$, namely abnormality prior network, to model the sample-wise abnormality prior based on $\textbf{p}_{x}$. This prior $\Omega_{x}=\psi(\textbf{p}_{x})$ is then incorporated as data-dependent abnormal features into the learnable abnormal tokens of the abnormality prompts in CAP to dynamically adapt the learned fine-grained abnormalities to a given target dataset, with each individual abnormality prompt refined as follows:
$\mathcal{\hat{P}}^{a_{i}}=[V_{1}][V_{2}]...[V_{E}][A_{1}^{i}\oplus\Omega_{x}][A_{2}^{i}\oplus\Omega_{x}]...[A_{E^{\prime}}^{i}\oplus\Omega_{x}][object],$ (4)
@@
The goal of DAP is to introduce sample-wise abnormality information. However, there is no abnormality from the top $M$ patches of normal images, and thus, simply applying the prior $\Omega_{x}$ to normal images would introduce noise into the learnable abnormal tokens, damaging the learning of fine-grained abnormalities. To address this issue, we propose an abnormality prior learning loss $\mathcal{L}_{prior}$ to enforce that $\Omega_{x}$ is the features mapped from the most abnormal $M$ patches if $x$ is an abnormal image, while it is minimized to be a null vector if it is a normal image. Formally, $\mathcal{L}_{prior}$ can be defined as follows:
@@ 3.5 Training and Inference
where I is a full-one matrix, $\mathcal{L}_{Focal}(\cdot)$ and $\mathcal{L}_{Dice}(\cdot)$ denote a focal loss (Lin et al., 2017) and a Dice loss (Li et al., 2019b), respectively. To ensure the accuracy of locating the top abnormal features in DAP, we apply the same learning objective to optimize the segmentation maps $\mathbf{\mathcal{M}^{n}}\in\mathbb{R}^{h\times w}$ and $\mathbf{\mathcal{M}^{a}}\in\mathbb{R}^{h\times w}$, which are derived from the normality-oriented scores ${S_{x}^{n}}$ and abnormality-oriented scores ${S_{x}^{a}}$, respectively.
@@
where $s_{a}^{\prime}(x)=\frac{1}{2}\left(\max({S_{x}^{a}})+\max({\hat{S}}_{x}^{a})\right)$ represents the average of the maximum anomaly scores from ${S_{x}^{a}}$ and ${\hat{S}_{x}^{a}}$. Following previous methods (Zhu & Pang, 2024; Chen et al., 2023a; Zhou et al., 2024; Jeong et al., 2023), $s_{a}^{\prime}(x)$ is treated as a complementary anomaly score to $s_{a}(x)$ and incorporated into Eq. 9, as $s_{a}^{\prime}(x)$ are helpful for detecting local abnormal regions. The image-level anomaly score $s(x)$ is then optimized by minimizing the following loss on $X_{train}$:
@@
$\mathcal{L}=\mathcal{L}_{local}+\mathcal{L}_{global}+\mathcal{L}_{prior}+\mathcal{L}_{oc},$ (11)
@@ 4 Experiments
Datasets. To verify the effectiveness of FAPrompt, we conduct extensive experiments across 19 publicly available datasets, including nine popular industrial defect inspection datasets on varying products/objects (MVTecAD (Bergmann et al., 2019), VisA (Zou et al., 2022), DAGM (Wieler & Hahn, 2007), DTD-Synthetic (Aota et al., 2023), AITEX (Silvestre-Blanes et al., 2019), SDD (Tabernik et al., 2020), BTAD (Mishra et al., 2021), MPDD (Jezek et al., 2021), and ELPV(Deitsch et al., 2019)) and ten medical anomaly detection datasets on different organs like brain, fundus, colon, skin and thyroid (BrainMRI (Salehi et al., 2021), HeadCT (Salehi et al., 2021), LAG (Li et al., 2019a), Br35H (Hamada, ), CVC-ColonDB (Tajbakhsh et al., 2015), CVC-ClinicDB (Bernal et al., 2015), Kvasir (Jha et al., 2020), Endo (Hicks et al., 2021), ISIC (Gutman et al., 2016), TN3K (Gong et al., 2021)) (see Appendix A for details about the datasets).
Table 1: Image-level ZSAD results (AUROC, AP) on 13 AD datasets. The best and second-best results are respectively highlighted in red and blue. The results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD are averaged performance across their multiple data subsets (see Appendix D for breakdown results).
@@
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Industrial MVTecAD (74.1, 87.6) (71.5, 86.4) (91.8, 96.5) (86.2, 93.5) (92.5, 96.7) (88.8, 94.8) (71.8, 84.9) (91.5, 96.2) (91.9, 95.7)
VisA (66.4, 71.4) (65.0, 70.2) (78.8, 81.4) (78.0, 81.4) (79.2, 81.7) (62.8, 68.1) (78.1, 82.3) (82.1, 85.4) (84.5, 86.8)
SDD (95.5, 87.9) (94.7, 77.9) (94.0, 87.2) (97.5, 93.4) (95.3, 91.3) (96.8, 90.0) (89.9, 50.4) (98.1, 93.4) (98.6, 95.9)
BTAD (34.5, 52.5) (51.0, 62.1) (68.2, 70.9) (73.6, 68.6) (80.3, 73.1) (66.8, 77.4) (48.4, 53.9) (88.3, 87.3) (92.0, 92.2)
MPDD (54.3, 65.4) (56.2, 66.0) (63.6, 69.9) (73.0, 80.2) (68.9, 71.9) (55.1, 64.2) (61.0, 69.1) (77.0, 82.0) (80.6, 83.3)
AITEX (71.0, 45.7) (71.5, 46.7) (73.0, 54.7) (57.6, 41.3) (72.5, 55.4) (66.2, 39.0) (48.6, 37.8) (62.2, 40.4) (71.9, 53.2)
DAGM (79.6, 59.0) (82.5, 63.7) (91.8, 79.5) (94.4, 83.8) (89.7, 76.3) (87.5, 74.6) (96.3, 85.5) (97.5, 92.3) (98.9, 95.7)
DTD-Synthetic (71.6, 85.7) (66.8, 83.2) (93.2, 92.6) (86.4, 95.0) (94.9, 97.3) (83.1, 91.9) (84.1, 92.9) (93.5, 97.0) (95.9, 98.3)
ELPV (59.2, 71.7) (69.4, 80.2) (74.0, 86.0) (65.5, 79.3) (70.6, 83.0) (73.0, 86.5) (78.4, 89.2) (81.5, 91.3) (83.5, 92.0)
Medical BrainMRI (73.9, 81.7) (80.6, 86.4) (86.6, 91.5) (89.3, 90.9) (88.7, 91.3) (61.3, 44.9) (78.2, 86.7) (90.3, 92.2) (95.5, 95.6)
HeadCT (56.5, 58.4) (60.0, 60.7) (81.8, 80.2) (89.1, 89.4) (81.6, 84.2) (78.4, 78.8) (80.3, 73.4) (93.4, 91.6) (94.8, 93.5)
LAG (58.7, 76.5) (58.2, 76.9) (59.2, 74.8) (73.6, 84.8) (65.1, 78.0) (69.6, 82.9) (72.6, 84.7) (74.3, 84.9) (75.6, 85.4)
Br35H (78.4, 78.8) (82.7, 81.3) (80.5, 82.2) (93.1, 92.9) (88.4, 88.9) (86.0, 87.5) (85.7, 89.1) (94.6, 94.7) (97.8, 97.5)
To assess the ZSAD performance, the models are trained on the MVTecAD dataset by default and evaluated on the test sets of other datasets without any further training or fine-tuning. We obtain the ZSAD results on MVTecAD by changing the training data to the VisA dataset.
Competing Methods and Evaluation Metrics. We compare our method, FAPrompt, with several state-of-the-art (SotA) methods, including five handcrafted text prompt-based methods – raw CLIP (Radford et al., 2021), CLIP-AC and WinCLIP (Jeong et al., 2023), APRIL-GAN (Chen et al., 2023a), and AnoVL (Deng et al., 2023) – and three learnable text prompt-based methods – CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), and AnomalyCLIP (Zhou et al., 2024). As for evaluation metrics, we follow previous works (Jeong et al., 2023; Zhou et al., 2024) and use two popular metrics: AUROC (Area Under the Receiver Operating Characteristic) and average precision (AP) to assess the image-level AD performance; for pixel-level AD performance, we employ AUROC and Area under per region overlap (PRO) to provide a more detailed analysis.
Implementation Details. Following previous approaches (Zhou et al., 2024; Chen et al., 2023a), we implement FAPrompt using the same CLIP implementation, OpenCLIP (Ilharco et al., 2021), using the publicly available pre-trained VIT-L/14@336px backbone. The parameters of both the visual and text encoders in CLIP are kept frozen. By default, learnable token embeddings are attached to the first nine layers of the text encoder, with a token length of four for each layer. The lengths of the learnable normal and abnormal text prompts are respectively set to five and two by default. The number of fine-grained abnormality prompts and selected patch tokens in the DAP module are both set to ten. We use the Adam optimizer with an initial learning rate of 1e-3. Further implementation details for FAPrompt and the competing methods are provided in Appendix B.
@@ 4.1 Main Results
Image-level ZSAD Performance. Table 1 presents the image-level ZSAD results of FAPrompt, compared to eight SotA methods across 13 AD datasets, including nine industrial defect AD datasets and four medical AD datasets. The results show that FAPrompt significantly outperforms the SotA models across almost all datasets. On average, compared to the best competing methods, it achieves up to 3.7% AUROC and 4.9% AP on industrial AD datasets and 5.2% AUROC and 3.4% AP on medical AD datasets. In particular, the weak performance of CLIP and CLIP-AC can be attributed to its over-simplified text prompt design. By utilizing more carefully designed handcrafted prompts, WinCLIP achieves better results than CLIP and CLIP-AC while preserving the training-free nature. APRIL-GAN and AnoVL improve over WinCLIP by using additional learnable layers and/or domain-aware tokens within the textual prompts. However, they heavily rely on sensitive handcrafted textual prompts and capture mainly coarse-grained semantics of abnormality, leading to poor performance when faced with anomalies that does not fit well to the pre-defined text descriptions, e.g., BTAD, MPDD, BrainMRI, HeadCT, and Br35H.
@@
As for text prompt learning methods, CoOp and CoCoOp are designed for general vision tasks, i.e., discriminating different objects, so they have weak capability in capturing the differences between normality and abnormality on the same object. AnomalyCLIP significantly improves performance by learning object-agnostic textual prompts for AD, demonstrating strong generalization capabilities across diverse datasets. However, AnomalyCLIP overlooks fine-grained abnormality details. FAPrompt overcomes this limitation via its two novel modules, CAP and DAP.
Table 2: Pixel-level ZSAD results (AUROC, PRO) on 14 AD datasets. The best and second-best results are respectively highlighted in red and blue. Note that medical datasets in Table 1 do not have pixel-level ground truth. Thus, different medical datasets are used here. Detailed breakdown results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD can be found in Appendix D.
Data Type Dataset Handcrafted Text Prompts Learnable Text Prompts
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp AnomalyCLIP FAPrompt
Industrial MVTecAD (38.4, 11.3) (38.2, 11.6) (85.1, 64.6) (87.6, 44.0) (89.8, 76.2) (33.3, 6.6) (86.7, 79.6) (91.1, 81.4) (90.6, 83.3)
VisA (46.6, 14.8) (47.8, 17.2) (79.6, 56.8) (94.2, 86.8) (89.9, 71.2) (24.1, 3.8) (93.6, 86.7) (95.5, 87.0) (95.9, 87.5)
SDD (28.4, 5.1) (33.5, 7.6) (95.9, 78.4) (93.0, 84.6) (97.9, 82.6) (91.8, 81.7) (93.7, 85.0) (98.1, 95.2) (98.3, 93.6)
BTAD (30.6, 4.4) (32.8, 8.3) (72.7, 27.3) (60.8, 25.0) (93.2, 62.8) (28.6, 3.8) (86.1, 72.0) (94.2, 74.8) (95.6, 75.2)
MPDD (62.1, 33.0) (58.7, 29.1) (76.4, 48.9) (94.1, 83.2) (84.0, 61.0) (15.4, 2.3) (95.2, 84.2) (96.5, 87.0) (96.5, 87.9)
AITEX (53.2, 15.3) (47.3, 11.8) (62.5, 41.5) (78.2, 68.8) (59.2, 49.1) (67.7, 54.9) (52.1, 56,9) (83.0, 66.5) (82.0, 62.6)
DAGM (28.2, 2.9) (32.7, 4.8) (87.6, 65.7) (82.4, 66.2) (92.0, 78.8) (17.5, 2.1) (82.8, 75.1) (95.6, 91.0) (98.3, 95.4)
DTD-Synthetic (33.9, 12.5) (23.7, 5.5) (83.9, 57.8) (95.3, 86.9) (97.5, 90.4) (55.8, 36.0) (93.7, 83.7) (97.9, 92.3) (98.3, 93.1)
Medical CVC-ColonDB (49.5, 15.8) (49.5, 11.5) (70.3, 32.5) (78.4, 64.6) (77.9, 49.8) (40.5, 2.6) (79.1, 69.7) (81.9, 71.3) (84.6, 74.7)
CVC-ClinicDB (47.5, 18.9) (48.5, 12.6) (51.2, 13.8) (80.5, 60.7) (82.1, 55.0) (34.8, 2.4) (83.4, 68.8) (82.9, 67.8) (84.7, 70.1)
Kvasir (44.6, 17.7) (45.0, 16.8) (69.7, 24.5) (75.0, 36.2) (72.5, 28.2) (44.1, 3.5) (79.1, 38.6) (78.9, 45.6) (81.2, 47.8)
Endo (45.2, 15.9) (46.6, 12.6) (68.2, 28.3) (81.9, 54.9) (80.5, 47.7) (40.6, 3.9) (83.1, 59.0) (84.1, 63.6) (86.4, 67.2)
ISIC (33.1, 5.8) (36.0, 7.7) (83.3, 55.1) (89.4, 77.2) (90.6, 79.8) (51.7, 15.9) (81.9, 68.9) (89.7, 78.4) (90.9, 81.2)
TN3K (42.3, 7.3) (35.6, 5.2) (70.7, 39.8) (73.6, 37.8) (80.9, 50.5) (34.0, 9.5) (72.4, 41.0) (81.5, 50.4) (84.5, 54.1)

### v2

@@ 3.4 Data-dependent Abnormality Prior Learning
One issue in ZSAD is that the fine-grained abnormality patterns can be very different from the auxiliary dataset to test datasets. In addition to the learning of a set of complementary fine-grained abnormality prompts, it is important to ensure that the learned fine-grained abnormality patterns are generalized to various test datasets. Inspired by the instance-conditional information design in CoCoOp [64], we introduce the DAP module to enhance the cross-dataset generalizability of the abnormal tokens in CAP by adaptively selecting the embeddings of the most abnormal regions to serve as a sample-wise abnormality prior for each image input. Particularly, given a query/test image $x$, DAP selects the most abnormal image patches as the abnormality prior to be fed into CAP for assisting the abnormality prompt learning. It achieves this by picking the top $M$ patches whose token embeddings are most similar to the abnormality prompt prototype $\mathbf{F}_{a}$:
@@
[…]
Let $\textbf{p}_{x}=\{p_{1},p_{2},...p_{M}\}$ be the top $M$ patch embeddings of $x$, FAPrompt then adds additional learnable layers $\psi(\cdot)$, namely abnormality prior network, to learn the sample-wise abnormality prior based on $\textbf{p}_{x}$. This prior $\Omega_{x}=\psi(\textbf{p}_{x})$ is then incorporated as data-dependent abnormal features into the learnable abnormal tokens of the abnormality prompts in CAP to dynamically adapt the fine-grained abnormalities to a given target dataset, with each individual abnormality prompt refined as follows:
$\mathcal{\hat{P}}^{a_{i}}=[V_{1}]...[V_{E}][A_{1}^{i}\oplus\Omega_{x}]...[A_{E^{\prime}}^{i}\oplus\Omega_{x}][object],$ (4)
@@
The goal of DAP is to introduce sample-wise abnormality information. However, there is no abnormality information from the top $M$ patches of normal images, and thus, simply applying the prior $\Omega_{x}$ to normal images would introduce noise into the learnable abnormal tokens, damaging the learning of fine-grained abnormalities. To address this issue, we propose an abnormality prior learning loss $\mathcal{L}_{prior}$ to enforce that $\Omega_{x}$ is the features mapped from the most abnormal $M$ patches if $x$ is an abnormal image, while it is minimized to be a null vector if it is a normal image. Formally, $\mathcal{L}_{prior}$ can be defined as follows:
@@ 3.5 Training and Inference
where I is a full-one matrix, $\mathcal{L}_{Focal}(\cdot)$ and $\mathcal{L}_{Dice}(\cdot)$ denote a focal loss [33] and a dice loss [31], respectively. To ensure the accuracy of locating the top abnormal features in DAP, we apply the same learning objective to optimize the segmentation maps $\mathbf{\mathcal{M}^{n}}\in\mathbb{R}^{h\times w}$ and $\mathbf{\mathcal{M}^{a}}\in\mathbb{R}^{h\times w}$, which are derived from the normality-oriented scores $S_{x}^{n}$} and abnormality-oriented scores $S_{x}^{a}$}, respectively.
@@
where $s_{a}^{\prime}(x)=\frac{1}{2}\left(\max({S_{x}^{a}})+\max({\hat{S}}_{x}^{a})\right)$ represents the average of the maximum anomaly scores from $S_{x}^{a}$} and $\hat{S}_{x}^{a}$}. Following previous methods [67, 10, 66, 25], $s_{a}^{\prime}(x)$ is treated as a complementary anomaly score to $s_{a}(x)$ and incorporated into Eq. 9, as $s_{a}^{\prime}(x)$ are helpful for detecting local abnormal regions. The image-level anomaly score $s(x)$ is then optimized by minimizing the following loss on $X_{train}$:
@@
Data Type Dataset Handcrafted Text Prompts Learnable Text Prompts
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp BLIP AnomalyCLIP FiLO FAPrompt
Industrial MVTecAD (74.1, 87.6) (71.5, 86.4) (91.8, 96.5) (86.2, 93.5) (92.5, 96.7) (88.8, 94.8) (71.8, 84.9) (72.3, 86.7) (91.5, 96.2) (91.2, 96.0) (91.9, 95.7)
VisA (66.4, 71.4) (65.0, 70.2) (78.8, 81.4) (78.0, 81.4) (79.2, 81.7) (62.8, 68.1) (78.1, 82.3) (57.8, 64.6) (82.1, 84.6) (83.9, 87.3) (84.6, 86.8)
BTAD (34.5, 52.5) (51.0, 62.1) (68.2, 70.9) (73.6, 68.6) (80.3, 73.1) (66.8, 77.4) (48.4, 53.9) (85.3, 91.2) (88.3, 87.3) (85.4, 91.5) (92.2, 92.5)
AITEX (71.0, 45.7) (71.5, 46.7) (73.0, 54.7) (57.6, 41.3) (72.5, 55.4) (66.2, 39.0) (48.6, 37.8) (71.3, 52.2) (62.2, 40.4) (78.2, 59.9) (74.1, 55.5)
DAGM (79.6, 59.0) (82.5, 63.7) (91.8, 79.5) (94.4, 83.8) (89.7, 76.3) (87.5, 74.6) (96.3, 85.5) (84.3, 66.6) (97.5, 92.3) (96.6, 90.4) (98.8, 95.3)
DTD-Synthetic (71.6, 85.7) (66.8, 83.2) (93.2, 92.6) (86.4, 95.0) (94.9, 97.3) (83.1, 91.9) (84.1, 92.9) (91.5, 96.3) (93.5, 97.0) (94.7, 98.0) (96.2, 98.1)
ELPV (59.2, 71.7) (69.4, 80.2) (74.0, 86.0) (65.5, 79.3) (70.6, 83.0) (73.0, 86.5) (78.4, 89.2) (75.2, 87.3) (81.5, 91.3) (82.2, 91.2) (83.7, 92.1)
SDD (95.5, 87.9) (94.7, 77.9) (94.0, 87.2) (97.5, 93.4) (95.3, 91.3) (96.8, 90.0) (89.9, 50.4) (85.8, 56.9) (98.1, 93.4) (97.9, 94.8) (98.4, 95.6)
MPDD (54.3, 65.4) (56.2, 66.0) (63.6, 69.9) (73.0, 80.2) (68.9, 71.9) (55.1, 64.2) (61.0, 69.1) (59.5, 67.7) (77.0, 82.0) (74.4, 76.9) (80.1, 83.9)
Medical BrainMRI (73.9, 81.7) (80.6, 86.4) (86.6, 91.5) (89.3, 90.9) (88.7, 91.3) (61.3, 44.9) (78.2, 86.7) (75.7, 83.0) (90.3, 92.2) (94.5, 94.9) (95.8, 96.2)
HeadCT (56.5, 58.4) (60.0, 60.7) (81.8, 80.2) (89.1, 89.4) (81.6, 84.2) (78.4, 78.8) (80.3, 73.4) (68.1, 62.8) (93.4, 91.6) (93.9, 92.1) (94.0, 92.4)
LAG (58.7, 76.5) (58.2, 76.9) (59.2, 74.8) (73.6, 84.8) (65.1, 78.0) (69.6, 82.9) (72.6, 84.7) (48.3, 66.8) (74.3, 84.9) (71.3, 83.8) (76.6, 86.1)
Br35H (78.4, 78.8) (82.7, 81.3) (80.5, 82.2) (93.1, 92.9) (88.4, 88.9) (86.0, 87.5) (85.7, 89.1) (76.4, 74.4) (94.6, 94.7) (97.7, 96.8) (97.6, 97.1)
Table 1: Image-level ZSAD results (AUROC, AP) on 13 AD datasets. The best and second-best results are respectively highlighted in red and blue. The results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD are averaged performance across their multiple data subsets (see Appendix D.1 for breakdown results).
@@ 4 Experiments
Datasets. To verify the effectiveness of FAPrompt, we conduct extensive experiments across 19 publicly available datasets, including nine popular industrial defect inspection datasets on varying products/objects (MVTecAD [4], VisA [69], DAGM [50], DTD-Synthetic [2], AITEX [44], SDD [45], BTAD [35], MPDD [26], and ELPV[15]) and ten medical anomaly detection datasets on different organs like brain, fundus, colon, skin and thyroid (BrainMRI [42], HeadCT [42], LAG [30], Br35H [22], CVC-ColonDB [46], CVC-ClinicDB [6], Kvasir [27], Endo [23], ISIC [21], TN3K [18]). To assess the ZSAD performance, the models are trained on the MVTecAD dataset by default and evaluated on the test sets of other datasets without any further training or fine-tuning. We obtain the ZSAD results on MVTecAD by changing the training data to the VisA dataset (see Appendix A for details about the datasets).
Competing Methods and Evaluation Metrics. We compare our FAPrompt with several state-of-the-art (SotA) methods, including five handcrafted text prompt-based methods (raw CLIP [39], CLIP-AC, WinCLIP [25], APRIL-GAN [10], and AnoVL [17]) and five learnable text prompt-based methods (CoOp [65], CoCoOp [64], AnomalyCLIP [66]), FiLO [20], and BLIP [29]). As for evaluation metrics, we follow previous works [25, 66] and use two popular metrics: Area Under the Receiver Operating Characteristic (AUROC) and Average Precision (AP) to assess the image-level AD performance; for pixel-level AD performance, we employ AUROC and Area Under Per Region Overlap (PRO) to provide a more detailed analysis. We also provide model complexity analysis and detailed qualitative results in Appendix C.1 and C.5, respectively.
Implementation Details. The implementation details for FAPrompt and the competing methods are provided in Appendix B.
@@
CLIP CLIP-AC WinCLIP APRIL-GAN AnoVL CoOp CoCoOp BLIP AnomalyCLIP FiLO FAPrompt
Industrial MVTecAD (38.4, 11.3) (38.2, 11.6) (85.1, 64.6) (87.6, 44.0) (89.8, 76.2) (33.3, 6.6) (86.7, 79.6) (85.1, 68.2) (91.1, 81.4) (92.3, 61.7) (90.6, 83.3)
VisA (46.6, 14.8) (47.8, 17.2) (79.6, 56.8) (94.2, 86.8) (89.9, 71.2) (24.1, 3.8) (93.6, 86.7) (85.7, 62.4) (95.5, 87.0) (95.9, 85.4) (95.9, 87.7)
BTAD (30.6, 4.4) (32.8, 8.3) (72.7, 27.3) (60.8, 25.0) (93.2, 62.8) (28.6, 3.8) (86.1, 72.0) (89.3, 60.7) (94.2, 74.8) (94.5, 56.3) (95.6, 75.1)
AITEX (53.2, 15.3) (47.3, 11.8) (62.5, 41.5) (78.2, 68.8) (59.2, 49.1) (67.7, 54.9) (52.1, 56,9) (59.0, 57.7) (83.0, 66.5) (70.7, 25.3) (82.0, 66.2)
DAGM (28.2, 2.9) (32.7, 4.8) (87.6, 65.7) (82.4, 66.2) (92.0, 78.8) (17.5, 2.1) (82.8, 75.1) (90.0, 78.0) (95.6, 91.0) (96.8, 90.6) (98.2, 95.0)
DTD-Synthetic (33.9, 12.5) (23.7, 5.5) (83.9, 57.8) (95.3, 86.9) (97.5, 90.4) (55.8, 36.0) (93.7, 83.7) (96.3, 88.2) (97.9, 92.3) (98.1, 88.6) (98.3, 93.3)
SDD (28.4, 5.1) (33.5, 7.6) (95.9, 78.4) (93.0, 84.6) (97.9, 82.6) (91.8, 81.7) (93.7, 85.0) (95.6, 76.4) (98.1, 95.2) (99.0, 82.9) (98.3, 94.1)
MPDD (62.1, 33.0) (58.7, 29.1) (76.4, 48.9) (94.1, 83.2) (84.0, 61.0) (15.4, 2.3) (95.2, 84.2) (90.2, 69.4) (96.5, 87.0) (95.7, 84.7) (96.5, 87.9)
Medical CVC-ColonDB (49.5, 15.8) (49.5, 11.5) (70.3, 32.5) (78.4, 64.6) (77.9, 49.8) (40.5, 2.6) (79.1, 69.7) (83.6, 66.9) (81.9, 71.3) (81.5, 63.9) (85.0, 73.3)
CVC-ClinicDB (47.5, 18.9) (48.5, 12.6) (51.2, 13.8) (80.5, 60.7) (82.1, 55.0) (34.8, 2.4) (83.4, 68.8) (79.5, 63.1) (82.9, 67.8) (84.6, 62.3) (84.7, 70.1)
Kvasir (44.6, 17.7) (45.0, 16.8) (69.7, 24.5) (75.0, 36.2) (72.5, 28.2) (44.1, 3.5) (79.1, 38.6) (76.5, 37.6) (78.9, 45.6) (85.0, 53.2) (82.1, 49.9)
Endo (45.2, 15.9) (46.6, 12.6) (68.2, 28.3) (81.9, 54.9) (80.5, 47.7) (40.6, 3.9) (83.1, 59.0) (78.5, 50.0) (84.1, 63.6) (87.1, 65.6) (86.8, 67.6)
ISIC (33.1, 5.8) (36.0, 7.7) (83.3, 55.1) (89.4, 77.2) (90.6, 79.8) (51.7, 15.9) (81.9, 68.9) (79.8, 57.6) (89.7, 78.4) (91.1, 80.1) (91.1, 81.6)
TN3K (42.3, 7.3) (35.6, 5.2) (70.7, 39.8) (73.6, 37.8) (80.9, 50.5) (34.0, 9.5) (72.4, 41.0) (73.9, 44.3) (81.5, 50.4) (79.8, 48.5) (84.7, 54.6)
Table 2: Pixel-level ZSAD results (AUROC, PRO) on 14 AD datasets. The best and second-best results are respectively highlighted in red and blue. Note that medical datasets in Table 1 do not have pixel-level ground truth. Thus, different medical datasets are used here. Detailed breakdown results for MVTecAD, VisA, DAGM, DTD-Synthetic, BTAD, and MPDD can be found in Appendix D.1.
@@ 4.1 Main Results
Image-level ZSAD Performance. Table 1 presents the image-level ZSAD results of FAPrompt, compared to ten SotA methods across 13 AD datasets, including nine industrial and four medical AD datasets. The results show that FAPrompt significantly outperforms the SotA models across almost all datasets. On average, compared to the best competing methods, it achieves up to 3.9% AUROC and 3.0% AP on industrial AD datasets and 2.3% AUROC and 1.3% AP on medical AD datasets. In particular, the weak performance of CLIP and CLIP-AC can be attributed to its over-simplified text prompt design. By utilizing more carefully designed handcrafted prompts, WinCLIP achieves better results than CLIP and CLIP-AC while preserving the training-free nature. APRIL-GAN and AnoVL improve over WinCLIP by using additional learnable layers and/or domain-aware tokens within the textual prompts. However, they heavily rely on sensitive handcrafted textual prompts and capture mainly coarse-grained semantics of abnormality, leading to poor performance on anomalies outside their seen domains (e.g., BTAD, MPDD, BrainMRI, HeadCT, Br35H).
@@
As for learnable prompt methods, BLIP pretrained on a smaller scale of data, so it is less effective as in CLIP-related method. CoOp and CoCoOp are designed for general vision tasks, i.e., discriminating different objects, so they have weak capability in capturing the subtle differences between normality and abnormality on the same object. AnomalyCLIP and FiLO significantly improves ZSAD performance by learning textual prompts. However, their prompt design either fails to capture fine-grained abnormality details or relies heavily on external tools (LLMs and Grounding DINO). They also ignore crucial abnormality information that can be derived directly from the query input. FAPrompt overcomes these limitations via the novel CAP and DAP.